# Assessing the Limitations of Large Language Models in Clinical Fact Decomposition

### v1

@@ G.2 Fact Atomicity
G.3 Fact Precision and Fact Recall
@@
G.4 Entailment Model Validation
@@
Note Type Junk Partial Examples
Procedures 9.0% 6.0% ___, _________, I., *, (Over), $\mid$, _______
Appendix H Qualitative Evaluation
Expert reviewers assessed overall completeness of fact decompositions and the correctness, independence, and atomicity of individual facts. Detailed results for Llama3-8B and Gemini are shown in Figure 9 and Figure 10 respectively.
@@ Appendix H Qualitative Evaluation
**footnotetext: We did not have a comprehensive qualitative evaluation done for o1-mini due to time constraints.
Appendix I Cost
@@
We adopt the prompt from Min et al. (2023), and include two in-context examples of fact decompositions of clinical notes written by a medical expert, illustrated in figure 11 The LLM is instructed to output independent facts as a delimited string. This approach resulted in fewer parsing errors than other methods.
@@
We adopt the prompt from Xie et al. (2024), instructs the LLM to output a binary 0/1 indicator of entailment in JSON format (figure 12). This prompt was preferred over those requesting rationales or chain-of-thought reasoning due to its efficiency in minimizing output tokens, reducing inference times, and lowering computational costs, especially given the large number of premise-hypothesis pairs in the FactEHR dataset.

### v2

@@ I.2 Fact Atomicity
I.3 Fact Precision and Fact Recall
@@
I.4 Entailment Model Validation
@@
Appendix J Qualitative Evaluation
Expert reviewers assessed overall completeness of fact decompositions and the correctness, independence, and atomicity of individual facts. Detailed results for Llama3-8B and Gemini are shown in Figure 9 and Figure 10 respectively  [^5] We did not have a comprehensive qualitative evaluation done for o1-mini due to time constraints..
@@
Appendix K Cost
@@
We adopt the prompt from (Min et al., 2023), and include two in-context examples of fact decompositions of clinical notes written by a medical expert, illustrated in figure 11. The LLM is instructed to output independent facts as a delimited string. This approach resulted in fewer parsing errors than other methods.
@@
We adopt the prompt from (Xie et al., 2024), instructs the LLM to output a binary 0/1 indicator of entailment in JSON format, as illustrated in figure 12. This prompt was preferred over those requesting rationales or chain-of-thought reasoning due to its efficiency in minimizing output tokens, reducing inference times, and lowering computational costs, especially given the large number of premise-hypothesis pairs in the FactEHR dataset.